OpenStudy (anonymous):

what is the gcf of 12 and 15 and how do you do it

OpenStudy (anonymous):

It is three remember what I told you: 12/15 divide by the 3 you get 4/5 which is the simplest form, so 3 is the GCF

OpenStudy (anonymous):

i did that but 4/5 cant = 3 lol can it

OpenStudy (anonymous):

I know but thats it you DONT divide it. Just simplify the fraction. What number did you divide by in the numerator and denominator to get it to the simplest form.

OpenStudy (anonymous):

No dont divide those two number. 12 /3 ___ 15 /3 The 3 is the common factor, the 3 the GREATEST common factor, because you can't divide it by anything else that is larger in order for it to be wholes number, or without getting a decimal in the numerator or denominator. Get it?

OpenStudy (anonymous):

sorry i dont thank you for your time >>>>

OpenStudy (zarkon):

12=3*2*2 15=3*5 what they have in common is a 3 so that is the answer

OpenStudy (anonymous):

Here are some videos that will help: http://www.youtube.com/watch?v=3eyjE-FN4Tw http://www.youtube.com/watch?v=3jGAcpK2V7Q